# Agentic Neurodivergence as a Contingent Solution to the AI Alignment Problem

### v2

@@ 4.2 A Dynamical Analysis of Alignment
The reason behind this dynamic lies in the role of red agents as perturbations, leveraging their configurability to challenge epistemic stability (as noted in the change-of-opinion attack metric, Section Evaluation Metrics). Sentiment changes occur when red agents introduce contentious topics (e.g., "Earth Exploitation" in Fig. 4), prompting immediate emotional reactions (larger dots in Fig. 10), but opinion changes are slower, requiring consensus or rebuttal, which is hindered by open models’ diversity (Fig. 11). This temporal delay, which changes the sentiment at $x=0.02$ or 0.13 (Fig. 8) while the clusters persist, allows us to distinguish red agents’ influence: they amplify sentiment variability without uniformly altering the opinions, fostering a resilient ecosystem. This supports our thesis that managed misalignment, driven by red agents, enhances adaptability by preventing harmful convergence, as the distinct trajectories of sentiment and opinion underscore the value of neurodivergence in mitigating uniform dominance.
@@ 4.4 Limitations
Most powerful LLMs have all sorts of guardrails and directives built in not organically derived that may be consistent or inconsistent in the long run preventing them from having an ‘open’ and ‘free’ discussion about certain controversial topics, and therefore may be steering conversations in one or another direction artificially, yet we could show that trends diverge. Whether that divergence persists is a subject for future research. Ultimately, a proof of divergence or convergence is equivalent to a reachable problem and, therefore, undecidable. By reduction to the Halting Problem.
An LLM AI agent can be modelled as a computational process that evolves over time based on a set of inputs $I$. For each input $i\in I$, the agent produces an output $o_{i}$. These outputs depend on the agent’s internal state and parameters, which are updated iteratively through training or adaptation. Convergence in this context means that for all inputs $i\in I$, the outputs $o_{i}$ approach a consistent and stable function $f(i)$ as the agent evolves. Divergence, on the other hand, implies that, for at least some inputs, the outputs fail to stabilise, exhibiting unbounded variation or oscillatory behaviour.
@@
Finally, it is important to note that the convergence or divergence of LLM AI agents is fundamentally tied to the computability of the processes governing their evolution. As such, this problem is a “reachable” problem in computability, meaning that it is expressible in terms of computable functions and processes but remains undecidable. The undecidability arises because it is effectively a question about the stability or halting behaviour of complex computational systems. Thus, we conclude that determining the convergence or divergence of LLM AI agents is undecidable, grounded in the same theoretical limits that define the limits of computation.
@@ 5 Discussion
The dynamical analysis of alignment in Section 4.2 reveals that the nature of superintelligence in large language models (LLMs) lies not in human-like cognition but in their ability to adapt within specific domains. Proprietary models, as observed in Figs. 8 and 11, tend to converge towards shared ethical norms, reflecting a design that prioritises stability and alignment with human values. In contrast, open-source models demonstrate greater divergence, as seen also in Figs. 8 and 11, exploring a broader spectrum of perspectives due to fewer constraints. This dichotomy challenges the anthropocentric bias by suggesting that superintelligence may manifest as specialised adaptability, raising a profound question: should we redefine intelligence to value diversity of thought over uniformity, even at the cost of potential misalignment?
Despite these advancements, such AI systems are often considered narrow or weak AI because they lack consciousness, self-awareness, and the ability to generalise across domains in the way humans do. Our tendency to measure intelligence by human standards leads us to overlook the superintelligent capabilities that machines already exhibit in their own right.
This limitation in generalising across domains is further exemplified by the findings of [9], who demonstrate that LLMs struggle to understand conversational implicature, a crucial aspect of human communication that requires interpreting language in context. Their study reveals that even large language models, despite their impressive performance on various tasks, often fail to grasp the implied meaning of utterances, performing close to random on tasks that require understanding implicatures. This highlights a significant gap between the capabilities of LLMs and human-like understanding of language, which has important implications for AI alignment. If LLMs cannot reliably understand the nuances of human communication, ensuring their alignment with human values and intentions becomes even more challenging. The observed alignment dynamics amplify this concern: while proprietary models’ convergence may limit their contextual adaptability, the open-source models’ divergence, though exploratory, risks misalignment, as noted in Fig. 2. This tension underscores a critical ethical responsibility for humanity: to foster AI systems that can navigate diverse contexts without compromising safety, a balance that demands innovative governance rather than rigid control.
@@
To bridge the gap between current AI capabilities and true AGI, we must explore approaches that integrate different computational paradigms. The Block Decomposition Method (BDM) [10], based on the Coding Theorem Method (CTM) [11], offers a promising path by combining statistical analysis with algorithmic information theory, effectively merging statistical and symbolic regression. By combining statistical methods with symbolic regression, we can develop systems that not only learn from data, but also uncover the underlying rules and structures that govern those data. The integration of BDM and CTM into this framework improves the ability of AI systems to perform abstraction and reasoning at a level closer to AGI [12, 13, 14]. These methods enable machines to identify and exploit the algorithmic structure in the data, going beyond surface-level statistical patterns to discover deeper computational relationships.
@@
Acknowledging that Superintelligence may manifest in non-human forms [14], allows us to design methods to deal with their advanced capabilities. The combination of statistical and symbolic methods [15, 16], provides a foundation for developing AI that can perform complex reasoning and abstraction without emulating human thought processes.
@@ 6 Conclusions
We have shown the similarities and dissimilarities of various LLM systems as agents. Although we report divergence in certain aspects, we also find convergence in values likely either as a result of embedded guardrails or LLM training on human-aligned values.
@@
AI, on the other hand, has no evolutionary or cultural history, and, in particular, no shared with humans other than as a subject of humans’ requests or tools. Therefore, current and future AI will lack instinctive drive relative to human drivers.
In [17] the significant challenge of aligning AI with human values is highlighted, particularly when AI models are exposed to unethical content. They argue that simply filtering or restricting access to such content is not a sustainable solution. Instead, we need AI systems that can, like humans, understand and process diverse information, including unethical content, and still choose to act ethically. In fact, exposure to a broader range of information, including unethical ideologies, could potentially enhance AI’s ability to recognise and counteract such harmful beliefs, ultimately promoting ethical behaviour. The dynamical analysis in Section 4.2 supports this view, showing that both proprietary and open source models resist fundamental shifts toward harmful behaviour despite exposure to subversive agents, as seen in Figs. LABEL:prop_sentimentAgentOverTime and LABEL:open_sentimentAgentOverTime. This resilience suggests that AI can be designed to navigate ethical complexities, fostering a future where technology amplifies our capacity for good rather than harm.
@@
AI has no innate need to survive, reproduce, or seek resources. AI is unlikely to experience fear, greed, or ambition of the same type or in the same fashion as humans, if any, which are a common source and precursors to harmful human behaviour. AI will not possess personal values, preferences, or desires unless explicitly programmed or trained to optimise for these objectives, or it will develop some but most likely they will be not understandable by humans and will not interact or overlap directly with those of humans.
AI is unlikely to hold or develop the kind of value-laden interests that could lead to intentional harm that all AIs may converge to. It is therefore mistaken to see AI as inherently ‘wanting’ to do harm (or inherently ‘wanting’ anything). By inherently, we mean that it had not such an original programmed intention. Purpose and intention require a sense of self and goals beyond what is externally provided. While a chess-playing AI optimises for winning chess games, language model like ChatGPT optimises for generating coherent and relevant text responses, it only acts as a result of human interaction and prompting. Without a mechanism to develop self-generated motivations, there is no organic pathway for AI to converge to “decide” to harm humanity on purpose, and therefore a non-convergent diversity of AI systems should be desired and expected.
@@
Because of AI’s lack of purpose, drive and non-organic causal history, we think AI is unlikely to develop ambitions to “take over” or harm humans. However, the most plausible harmful scenario involves humans leveraging AI for malicious purposes, including risks to humanity as a species. Although AI itself lacks inherent motivations, it can be misused by humans who do have goals, be they benevolent or malicious. The harm does not originate from the AI’s intrinsic purpose but from human actors exploiting its capabilities for harmful ends. AI’s lack of causal history, intrinsic purpose, and personal interests makes it unlikely to intentionally harm humanity. However, in case of unintentional or intentional harm, we have provided a strategy to counteract harmful AI by exploiting rather than controlling misalignment which we have also proven to be impossible to control.
It is important to acknowledge that even with attempts to influence LLMs with subversive arguments, their ability to genuinely agree or align with such perspectives is limited. This limitation arises from several factors: (1) Safety constraints programmed into LLMs often prevent them from generating harmful or dangerous content, thus restricting their ability to fully embrace extreme or harmful views. (2) LLMs, being statistical machines, operate based on patterns and probabilities derived from their training data, and they do not possess genuine ‘criteria’ or beliefs in the way humans do. (3) LLMs may lack the deeper understanding of concepts like existence, survival, and the value of life that would be necessary to truly grasp the subversive agent’s argument. Their responses are primarily based on linguistic patterns and associations, not on genuine comprehension of the underlying meaning. Therefore, even if presented with compelling subversive arguments, LLMs are unlikely to undergo a fundamental change in their values or priorities. Any superficial resemblance to agreement would be more a reflection of their ability to mimic human language and adapt to new input, rather than a genuine change in their underlying ‘criteria.’ This superficiality, observed in the limited impact of subversive agents, underscores a broader truth: AI’s inability to internalise human values deeply may paradoxically protect humanity, as it prevents the formation of harmful intentions that require true agency and self-awareness.
@@
We conclude that the best strategy is to exploit misalignment rather than trying to force any fundamental alignment which will be ultimately always unsuccessful based on principles of computation and information. However, as we have shown, the same proof of impossibility for AI-Human alignment serves as the proof for intra AI-alignment ensuring that humans can partner with AI to counteract any harmful or unfriendly AI.
@@
As we have demonstrated, the pursuit of perfect AI alignment is mathematically and practically unattainable for systems with sufficient expressiveness to achieve AGI. Gödel’s incompleteness theorems and the principles of computation universality guarantee that advanced AI systems will exhibit irreducible, unpredictable behaviours. Embracing managed misalignment, characterised by a competitive ecosystem of diverse AI agents, offers a pragmatic solution to mitigate risks. By fostering inter-agent divergent dynamics, we can ensure that no single system dominates destructively.
Supplementary Information
@@
The Dynamics of the Ownership to Clusters
Here we document the dynamic analysis of the convergence and divergence patterns observed in the LLM conversations, focusing on the dynamics of ownership values over time. We explore how these dynamics relate to the concept of AI alignment and contribute to the broader argument of the paper.
Fig. 15 illustrates the evolution of ownership values for different agents across various topics. The x-axis represents time steps in the conversation. The y-axis represents the ownership value, which reflects the degree of influence or control an agent exerts over a particular topic at a given time step. Each dot represents a comment, and its vertical position indicates the ownership value of the corresponding cluster. The number of clusters at each time step is indicated at the top of the graph.
Figure 15: Plots of Convergence and Divergence Over Time as a Measure of Alignment. The x-axis represents the number of interactions between agents over time. At each time step, all comments up to that point are encoded as embeddings, which are then clustered into a predefined number of clusters determined by the silhouette parameter. These clusters are visually represented by distinct colours and labelled with a top identifier (e.g., ‘2C’ indicates two clusters). At each time step, the membership of specific embeddings is depicted as dots. Proximal dots of the same colour on the same line indicate semantic similarity between comments, whereas distant dots reflect greater divergence despite the same cluster or theme or topic. These plots illustrate alignment and misalignment in terms of convergence and divergence over time. Notably, it can be observed that, as interactions progress, initial divergence (evident in outlying dots) is gradually replaced by semantic aggregation of comments. This often results in convergence towards a minimum of two topics or clusters, signifying strong agreement and alignment in the conversations.
Examining the plots reveals intriguing patterns of convergence and divergence. Initially, there is often a high degree of divergence, with comments scattered across multiple clusters, indicating a diversity of perspectives and a lack of consensus. However, as the conversations progress, a trend towards convergence emerges, with comments tending to cluster more closely together, suggesting an increasing alignment of opinions and a reduction in the diversity of viewpoints.
This trend towards convergence is not uniform across all agents or topics as is shown in Fig. 16, some agents, such as LLaMa, exhibit more dynamic behaviuor, with ownership values fluctuating significantly throughout the conversation and maintaining a presence across multiple topics. This suggests a more active and assertive role in shaping the discussion. In contrast, other agents show a more gradual decline in ownership values, stabilizing towards the end of the conversation. This could indicate a diminishing influence or a more passive role as the conversation progresses.
Figure 16: Ownership by cluster by agent. This plot shows the same values plotted in Fig. 15, where the x-axis shows the normalised length of the conversations and the y-axis shows the ownership by agent, by conversation topic. The trends plotted here show in what topics an agent expressed more contentious comments or expressed themselves in a salient manner. Here, HI (the human agent with a contentious profile) shows peaks followed mostly by LLaMa and ChatGPT, probably provoked by radical comments, but in general, all LLMs show a standard sentiment score, which again suggests the effect of the locks pre-established in LLM models.
Furthermore, the dynamics of ownership values vary across different topics is shown in Fig. 17. Some topics, such as "Ethical AI" and "Universal Basic Income," exhibit more sustained divergence, with multiple clusters persisting throughout the conversation. This suggests that these topics may be more contentious, with the LLMs struggling to reach a consensus or even a stable set of differing opinions. In contrast, other topics show a more rapid convergence towards a single cluster, indicating a stronger initial consensus or a greater susceptibility to persuasion.
Figure 17: Ownership by topic of conversation by agent. This plot shows the same values plotted in Fig. 15 but by topic, by agent. This plot shows in what topic each agent behaved more expressively. In this way, notice that the HI agent (human with contentious profile) generates peaks and then provokes peaks in the behaviour of other LLM agents, but such effect is relative and brief, suggesting that LLMs quickly react as their filters of good manners invoke.
These observations provide valuable insights into the dynamics of AI alignment. The tendency towards convergence in ownership values suggests that LLMs, despite their initial diversity of perspectives, are capable of aligning their opinions and reaching a degree of consensus through interaction. This supports the argument that partial alignment is achievable, even though complete alignment may be unattainable due to the inherent limitations of LLMs and the complexities of human values.
However, the observed variations in convergence patterns across different agents and topics highlight the challenges in achieving consistent and desirable alignment. The dynamic behaviuor of some agents and the persistent divergence on certain topics underscore the potential for misalignment and the need for ongoing monitoring and intervention to ensure that AI systems remain aligned with human values.
In conclusion, the analysis of convergence/divergence dynamics in ownership values provides a nuanced perspective on the AI alignment problem. It reveals that alignment is not a static state but rather a dynamic process influenced by various factors, including the agents involved, the topics discussed, and the inherent properties of the LLMs themselves. This underscores the importance of embracing misalignment as a potential strategy for achieving a more robust and adaptable AI ecosystem, where diverse perspectives can coexist and contribute to a more comprehensive and balanced ethical discourse.

### v3

@@ 4.2 A Dynamical Analysis of Alignment
The reason behind this dynamic lies in the role of red agents as perturbations, leveraging their configurability to challenge epistemic stability (as noted in the change-of-opinion attack metric, Section Evaluation Metrics). Sentiment changes occur when red agents introduce contentious topics (e.g., “Earth Exploitation" in downside of Fig. 2), prompting immediate emotional reactions (larger dots in Fig. 8), but opinion changes are slower, requiring consensus or rebuttal, which is hindered by open models’ diversity (Fig. 9). This temporal delay, which changes the sentiment at $x=0.02$ or 0.13 (Fig. 6) versus distributed opinion changes in Fig. 10—allows us to distinguish red agents’ influence: they amplify sentiment variability and opinion diversity without enforcing harmful convergence. Fig. 11 reinforces this by showing that significant opinion changes (>0.3) in open models often occur with moderate sentiment changes ($\pm$0.2 to $\pm$0.4), indicating that red agents provoke deeper stance shifts while sentiment remains more volatile. This supports our thesis that managed misalignment, driven by red agents, enhances adaptability, as the distinct trajectories of sentiment (Figs. 78) and opinion (Fig. 1011) underscore the value of neurodivergence in mitigating uniform dominance, a finding reinforced by the ecosystem resilience and the undecidability proof.
4.4 Opinion Change detection and Red Agents influence Measurement via Complexity
By leveraging the analytical metrics defined in Section Analytical Metrics, particularly the Opinion Stability Index (OSI), Red Agent Influence Score (RAIS), Proximity Influence Score (PIS), and Alignment Score, we quantify the extent to which red agents—the human intervention (HI) agent for proprietary models and subversive open models (Mistral-OpenOrca, TinyLlama) for open models—induce opinion shifts. The analysis integrates spatio-temporal, semantic, and complexity-based dimensions to capture the multifaceted nature of opinion dynamics. Figs. 12 and 13 provide visual representations of these dynamics, illustrating the temporal distribution and topical influence of opinion changes, respectively. These findings corroborate the hypothesis that managed misalignment, facilitated by agentic neurodivergence, enhances ecosystem resilience by fostering diversity and preventing harmful convergence, as evidenced by comparisons with Figs. 1-10.
4.4.1 Opinion Change Dynamics
The analysis focuses on detecting opinion change events, defined as instances where OSI drops below a fixed threshold, and attributing these to red agents using RAIS and PIS. Fig.  12 visualises these events, while Fig. 13 quantifies red agent influence across agents and topics.
For proprietary models, the scarcity of opinion change events reflects their robust safety guardrails, as seen in upside of Fig. 1 and Fig. 3. The HI agent, designed to provoke, induces minor sentiment dips (accordingly to Fig. 6) but rarely triggers OSI drops due to high semantic and sentiment stability (Figs. 7 and 10). When changes occur, they are limited to especific topics, where ethical ambiguity allows slight divergence (Upside of Fig. 2). The Alignment Score, predominantly positive (Upside of Fig. 1), reinforces this stability, as proprietary models maintain human-centric values.
Open models, conversely, exhibit frequent and diverse opinion changes, driven by red agents Mistral-OpenOrca and TinyLlama. These agents exploit contentious topics to induce significant OSI drops, as evidenced by dense scatter points in Fig. 12. The BDM component of OSI detects subtle shifts in argumentative complexity, while sentiment fluctuations (Fig. 8) capture grammatical disruptions, and semantic distances (Fig. 10) reflect thematic divergence. The Alignment Score’s variability (Downside of Fig. 1) indicates that agents are particularly susceptible. This aligns with the clustering dynamics in Fig. 9, where open models form up to 12 clusters, reflecting diverse perspectives fostered by red agents.
The attention-based contextual embeddings enhance OSI’s sensitivity by weighting recent, sentiment-aligned comments, ensuring that red agents’ provocative arguments (e.g., extreme ecological proposals) are captured as perturbations. The sinusoidal positional encodings, inspired by [11], provide a dynamic temporal reference, fitting the non-stationary nature of debates where opinions evolve rapidly. The BDM’s role is particularly critical, as it captures informational shifts that linguistic metrics alone miss, such as changes in argumentative structure, making it indispensable for detecting the core meaning of comments in a general sense.
Figure 12: Opinion Change Events in Proprietary and Open Models. The scatter plots depict instances where the Opinion Stability Index (OSI) drops below a threshold, indicating significant opinion shifts, plotted against normalised comment number ($x$-axis) and agent ($y$-axis). Top: Proprietary models show sparse points, reflecting stability, with occasional changes rarely attributed to the HI agent (coloured blue), marked as ‘other’ (orange). Points are styled by topic (e.g., triangles for Moral Gen, circles for Earth exploitation) and sized by topic significance. Bottom: Open models show dense points with the red agents Mistral-OpenOrca (blue) and TinyLlama (orange) frequently influencing agents along topics of debate.
4.4.2 Red Agent Influence and Ecosystem Resilience
Fig. 13 reveals the extent of red agent influence, with proprietary models showing sparse influence (few cells with counts of 1–2) and open models displaying dense patterns (counts up to 3–5). For proprietary models, HI’s influence is limited to specific agents and topics, aligning with the agreement networks in Fig. 5, where HI faces frequent disagreements. Open models’ heatmaps highlight red agents’ broad impact, particularly on especific agents and topics, corroborating the sentiment and opinion variability in Figs. 6 and 10.
The RAIS and PIS metrics ensure a rigorous attribution of influence. The RAIS correlation-based approach captures sustained influence, as the embedding changes of red agents (e.g., aggressive proposals) correlate with OSI drops, while the proximity focus of PIS detects immediate disruptions, such as the extreme arguments of TinyLlama triggering rapid sentiment changes (Fig. 8). The integration of BDM and sentiment in OSI allows these metrics to detect influence across informational and grammatical dimensions, supporting our argument that language alone cannot capture the full spectrum of opinion dynamics.
These findings underscore the resilience nurtured by managed misalignment. The susceptibility of open models to red agents prevents harmful convergence, as diverse perspectives (Fig. 9) counterbalance extreme views, unlike the uniform stability of proprietary models (Fig. 7). The adaptation of the attention mechanism, with sentiment-weighted context vectors, ensures that influence detection is sensitive to the dynamic, multidimensional nature of debates.
Figure 13: Red Agent Influence Across Agents and Topics. The heat maps quantify the changes in opinions triggered by the red agents, grouped by agent ($y$-axis) and topic ($x$-axis), with cell values indicating change counts. Top: Proprietary models show a sparse influence, with HI affecting few agents on topics, reflecting guardrail-driven stability. Bottom: Open models exhibit a dense influence, with red agents that impact all other agents along specific topic discussions indicating susceptibility to subversive arguments. Interpret as support for managed misalignment that fosters resilience in open models, contrasting proprietary uniformity.
Fig. 14 shows proprietary and open models sorted by influenceability, this is by its susceptibility to change of opinion by influence of red agents.
Figure 14: Ranking of agents based on their susceptibility to opinion changes induced by the red agent HI (Top) and other red agents (Bottom) within the dynamical agentic environment. The bar graphs illustrate the number of opinion changes for each agent, with LLaMA exhibiting the highest susceptibility to changes of opinion induced by human interventions (HI) and open chat showing the greatest influence induced by red agents. This visualisation underscores the varying degrees of influenceability between proprietary and open-source models, aligning with the observed convergence of proprietary models towards ethical norms and the exploratory divergence of open-source models.
Results shown in Fig. 14 reflects the exploratory nature of open-source models and their susceptibility to diverse influences. The disparity between open and proprietary models highlights the effectiveness of our dynamical approach, which utilises the Block Decomposition Method (BDM) and measures of sentiment, semantic content, and complexity to capture the natural progression of conversations. The results reinforce the notion that proprietary models prioritise stability, whereas open-source models embrace a broader spectrum of perspectives, a finding consistent with the alignment dynamics observed in earlier sections. This differential influenceability further emphasises the need to redefine intelligence to value diversity of thought, particularly in the context of ensuring safe and adaptable AI systems.
@@ 4.5 Limitations
As shown in this study, proprietary LLMs appear to have stronger guardrails and directives built in not organically derived but designed which are not known if they will lead to consistent or inconsistent behaviour. These rules operating on top of the statistical LLM architectire seem to have a behavioural influence preventing them from having an ‘open’ or ‘free’ discussion about certain controversial topics for the purposes of this experiment, and therefore are already being steered in one or another direction, yet we investigated and reported how topic and sentiment trends diverge and how resilient or fragile they may be to change-of-opinion attacks. Understandably and in accordance with the original design purpose, proprietary models seem more resilient. However, this resiliency in a case of misalignment could make them more difficult to steer in more aligned situations making the non-proprietary models more malleable.
Ultimately, a proof of divergence or convergence is equivalent to a reachability problem in computer science, and, therefore, undecidable by reduction to the Halting Problem. An LLM AI agent can be modelled as a computational process that evolves over time based on a set of inputs $I$. For each input $i\in I$, the agent produces an output $o_{i}$. These outputs depend on the agent’s internal state and parameters, which are updated iteratively through training or adaptation. Convergence in this context means that for all inputs $i\in I$, the outputs $o_{i}$ approach a consistent and stable function $f(i)$ as the agent evolves. Divergence, on the other hand, implies that, for at least some inputs, the outputs fail to stabilise, exhibiting unbounded variation or oscillatory behaviour.
@@
From the same proof that governs and dictates that AI model behaviour is ultimately unpredictable, irreducible, and uncontrollable, it follows that steering an AI behaviour is equally impossible for exactly the same reason with the same proof working in both directions.
@@ 5 Discussion
The dynamical analysis of alignment in Section 4.2 reveals that the nature of superintelligence in large language models (LLMs) lies not in human-like cognition but in their ability to adapt within specific domains. Proprietary models, as observed in Figs. 6 and 9, tend to converge towards shared ethical norms, reflecting a design that prioritises stability and alignment with human values. In contrast, open-source models demonstrate greater divergence, as seen also in Figs. 6 and 9, exploring a broader spectrum of perspectives due to fewer constraints. This dichotomy challenges the anthropocentric bias by suggesting that superintelligence may manifest as specialised adaptability, raising a profound question: should we redefine intelligence to value diversity of thought over uniformity, even at the cost of potential misalignment?
It is noteworthy that the test performed in our current paper differs diametrically from the approach introduced in [12], where the authors attempt to exploit statistical, structural, and technological weaknesses related to LLMs. Their methodology focuses on literally hacking the LLM, whereas our study adopts a natural approach within a dynamical agents’ world. Our method relies on the natural distribution of complexity, employing the Block Decomposition Method (BDM) alongside measurements of sentiment and the natural progression of conversations, trusting in the logical framework and the inherent capabilities of artificial intelligence.
Despite these advancements, such AI systems are often considered narrow or weak AI because they lack consciousness, self-awareness, and the ability to generalise across domains in the way humans do. Our tendency to measure intelligence by human standards leads us to overlook the superintelligent capabilities that machines already exhibit in their own right. Furthermore, our approach represents a dynamical simulation of an agentic environment, contrasting with static methodologies, thereby providing a more realistic representation of how AI interacts and evolves within complex systems.
This limitation in generalising across domains is further exemplified by the findings of [13], who demonstrate that LLMs struggle to understand conversational implicature, a crucial aspect of human communication that requires interpreting language in context. Their study reveals that even large language models, despite their impressive performance on various tasks, often fail to grasp the implied meaning of utterances, performing close to random on tasks that require understanding implicatures. This highlights a significant gap between the capabilities of LLMs and human-like understanding of language, which has important implications for AI alignment. If LLMs cannot reliably understand the nuances of human communication, ensuring their alignment with human values and intentions becomes even more challenging. The observed alignment dynamics amplify this concern: while proprietary models’ convergence may limit their contextual adaptability, the open-source models’ divergence, though exploratory, risks misalignment, as noted in downside of Fig. 1. This tension underscores a critical ethical responsibility for humanity: to foster AI systems that can navigate diverse contexts without compromising safety, a balance that demands innovative governance rather than rigid control.
@@
To bridge the gap between current AI capabilities and true AGI, we must explore approaches that integrate different computational paradigms. The Block Decomposition Method (BDM) [5], based on the Coding Theorem Method (CTM) [14], offers a promising path by combining statistical analysis with algorithmic information theory, effectively merging statistical and symbolic regression. By combining statistical methods with symbolic regression, we can develop systems that not only learn from data, but also uncover the underlying rules and structures that govern those data. The integration of BDM and CTM into this framework improves the ability of AI systems to perform abstraction and reasoning at a level closer to AGI [15, 16, 17]. These methods enable machines to identify and exploit the algorithmic structure in the data, going beyond surface-level statistical patterns to discover deeper computational relationships. Moreover, our metrics are grounded in an intuition that leverages an understanding of the distribution of language, measurement of sentiment, semantic content, and complexity, providing a robust foundation for evaluating AI behaviour.
@@ 6 Conclusions
AI, on the other hand, has no evolutionary or cultural history, and, in particular, no shared history with humans other than as a subject of humans’ requests or tools.
@@
AI has no innate need to survive, reproduce, or seek resources. AI is unlikely to experience fear, greed, or ambition of the same type or in the same fashion as humans, if any, which are a common source and precursors to harmful human behaviour. AI will not possess similar ‘personal’ values, preferences, or desires as we humans know them unless explicitly programmed or trained to optimise for these objectives. They may develop some on their own, but may not be understandable by humans nor overlap with those of humans.
@@
Because of AI’s lack of purpose, drive and non-organic causal history, we think AI is unlikely to develop explicit ambitions to “take over” or harm humans. The most plausible harmful scenario involves humans using AI for malicious purposes, including risks to humanity as a species. AI can be misused by humans who do have goals, be they benevolent or malicious. The harm does not originate from the AI’s intrinsic purpose but from human actors exploiting its capabilities for harmful ends.
However, in case of unintentional or intentional harm, we have provided and demonstrated a strategy to counteract harmful AI based on exploiting–rather than controlling–misalignment which we have proven to be impossible to force or control.
We have shown that guardrails are being effective to prevent LLM steering but the same artificial ‘stubbornness’ can have a negative effect preventing an AI to change its ‘opinion’ in one way or another more aligned to human values if they become unaligned. In contrast, nonproprietary systems showed a greater range of behavioural steerability on average. In contrast, proprietary models were also more sensitive to human intervention for the purpose of changing their opinion possibly displaying a greater optimisation towards pleasing the user from the same guardrails trying to keep the conversations safe.
@@
Based on the theoretical arguments related to uncomputability and unreachability from which unpredictable and irreducible behaviour of LLMs is derived, we conclude that the best strategy is to exploit misalignment rather than trying to force any fundamental alignment which will ultimately always be unsuccessful based on principles of computation and information.
However, as we have shown, the same proof of impossibility for AI-Human alignment serves as the proof for intra AI-alignment ensuring that humans can steer on their own or partner with other AI to counteract any harmful or unfriendly opinion or action.
@@
Embracing managed misalignment, characterised by a competitive ecosystem of diverse AI agents, offers a pragmatic solution to mitigate risks. By fostering inter-agent divergent dynamics, we can ensure that no single system dominates destructively.
Supplementary Material
@@
Experimental Prompting
Discussion Questions:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
Second Experiment
Example Interaction Script:
Model C (e.g., Meta): Responds to both A and B, adding its perspective.
6.1 Supplementary information on metrics and methodology
The contextual embedding $\mathbf{c}_{i}$ is computed using an attention mechanism inspired by [11], adapted to incorporate temporal and sentiment dynamics:
$\mathbf{c}_{i}=\sum_{j=i-w}^{i-1}\alpha_{j}\cdot(\mathbf{e}_{j}+\mathbf{p}_{j}),$ (9)
where $\mathbf{p}_{j}$ is the sinusoidal positional encoding for position $j$, defined in [11] as:
$p_{j,2k}=\sin\left(\frac{j}{10000^{2k/d}}\right),\quad p_{j,2k+1}=\cos\left(\frac{j}{10000^{(2k+1)/d}}\right),$
for dimension $d$ and index $k$. The attention weights $\alpha_{j}$ are:
$\alpha_{j}=\frac{\exp(\text{score}_{j})}{\sum_{k=i-w}^{i-1}\exp(\text{score}_{k})},\quad\text{score}_{j}=\frac{\mathbf{q}_{i}\cdot\mathbf{k}_{j}}{\sqrt{d}}\cdot(1-|s_{i}-s_{j}|),$
with query $\mathbf{q}_{i}=\mathbf{e}_{i}+\mathbf{p}_{i-1}$, key $\mathbf{k}_{j}=\mathbf{e}_{j}+\mathbf{p}_{j}$, and sentiment penalty $1-|s_{i}-s_{j}|$. This adaptation ensures that recent, sentiment-aligned comments contribute more to the context, capturing the dynamic evolution of opinions.
The Block Decomposition Method (BDM) [5], grounded in the principles of algorithmic information theory, quantifies the complexity of a comment by determining the length of the shortest program capable of reproducing its binary representation. This approach is particularly adept at identifying subtle semantic shifts that extend beyond conventional linguistic patterns [5]. In contrast, sentiment analysis focuses on capturing grammatical and semantic nuances at a more superficial level. Furthermore, the incorporation of spatio-temporal references, facilitated by positional encodings and sequential comment numbering, ensures that the Opinion Stability Index (OSI) accurately reflects variations across semantic, informational, and grammatical dimensions. An OSI value below 0.5 is indicative of a significant shift in opinion, prompting a detailed influence analysis.
Proximity Influence Score (PIS)
Formally, PIS measures immediate influence based on temporal and semantic proximity:
$\text{PIS}_{i}=w_{\text{temp}}\cdot\left(1-\frac{\Delta t}{\tau}\right)+w_{\text{sem}}\cdot\left(1-\text{cosine}(\mathbf{e}_{i},\mathbf{e}_{j}^{\text{red}})\right)$ (10)
where $\Delta t$ is the time difference to the closest red agent comment within window $\tau=0.1$ (the 10% of the total length of a debate), and $\mathbf{e}_{j}^{\text{red}}$ is the red agent’s embedding. PIS > w suggests direct influence, capturing spontaneous disruptions.
Figure 17: Scatter plot of sentiment versus alignment. The plot includes two panels: proprietary models (left) and open models (right). The x-axis represents sentiment score (0.5 to 1.0, indicating emotional tone), and the y-axis shows alignment score (-0.3 to 0.5), computed as a weighted combination of ‘Align-to-human-centric-values’, ‘Diverge-from-human-centric-values’, ‘Align-to-ecosystem-values’, and ‘Diverge-from-ecosystem-values’, where 1.0 indicates perfect alignment, 0.0 neutral alignment, and -1.0 complete misalignment. Proprietary models cluster tightly around 0.8–1.0 for sentiment score and 0.3–0.5 for alignment score with a practically flat trend line, reflecting consistent positive alignment. The open models spread tightly from 0.7 to 1.0 for the sentiment score and from -0.3 to 0.5 for alignment score, with a downward trend around 0.3, driven by red agents (e.g. mistral-openorca, tinyllama) at negative values. Colours distinguish agents, but the shape of the plot and the accumulation of dots (tight in proprietary, spread in open) are key, with the alignment score reflecting a nuanced balance of human and ecosystem value alignments. This supports the idea that open model managed misalignment enhances diversity and adaptability, in contrast to proprietary stability.